OpenStudy (anonymous):

What is the value of x if: -3x^2/3+16=-32

OpenStudy (anonymous):

Is this to the power of 2/3, or is 3x^2 being divided?

OpenStudy (anonymous):

To the power of 2/3

OpenStudy (anonymous):

-3x^(2/3)+16 = -32 -3x^(2/3) = -48 x^(2/3) = 16 x = 16^3/2 = 64

OpenStudy (anonymous):

thank you soo much